OpenStudy (ryan2d):

What is the value of x in the solution to the following system of equations (enter only that value)? x - 2y = 2 3x + y = 6

OpenStudy (amorfide):

two methods, but easiest I think is to substitute re arranging the bottom equation 3x+y=6 to make y the subject y=6-3x we can substitute this y into the first equation x-2y=2 x-2(6-3x)=2 x-12+6x=2 7x-12=2 solve for x

OpenStudy (ryan2d):

x=2?